# Rank three representations of Painlevé systems: I. Wild character varieties

### v1

@@ 1 Introduction
Since their discovery at the beginning of the 20th century, the Painlevé equations have attracted a lot of attention, and found a plethora of applications ranging from the theory of random matrices to integrable systems and quantum field theory. A classical result in the field due to Fuchs and Garnier (see Jimbo and Miwa [12, Appendix C]) is that the Painlevé equations all admit descriptions as isomonodromy equations of rank $2$ connections with (regular and irregular) singularities on the Riemann sphere. Meanwhile, Flaschka and Newell [9] found a different such rank $2$ isomonodromy equation description of the second Painlevé equation Painlevé II. This development pointed out that the Painlevé equations may admit several isomonodromic interpretations. Such alternative interpretations (known in the literature as Lax pair representations) are important in order to recognize the relevance of the Painlevé equations for further potential applications. Another Lax pair representation, this time in terms of rank $3$ singular connections, for the Painlevé VI equation was found by Harnad [10, Section 3c]. Subsequently, it was proved by Mazzocco [18] that Harnad’s duality can in fact be considered as an instance of Fourier–Laplace transformation of $\mathcal{D}$-modules. Later, Joshi, Kitaev and Treharne [13][14] gave a list of rank $3$ Lax (or, using their terminology, Fuchs–Garnier) pair representations of all the Painlevé equations. [^3] We note, however, that for Painlevé III they did not quite find a connection, because the differential term is multiplied by a singular matrix.
@@
Character varieties of (irregular) surfaces are of interest and admit applications throughout mathematics and mathematical physics. There are also quite interesting recent conjectures about them, for instance the Geometric P=W conjecture. Even though many theoretical results are known in the irregular case, relatively few concrete examples exist in the literature, especially in the higher (than $2$) rank case. This study illustrates how the general theory works by providing a detailed description of a family of rank $3$ examples that is known to be related to the well-known Painlevé family.
@@ 3 Joshi-Kitaev-Treharne systems
in some holomorphic local trivialization of $E$ around each $c_{i}$, where $Q$ is the irregular type from (3), $\Lambda$ is some constant $3\times 3$ matrix (called the residue matrix), and $O(1)$ refers to the holomorphic part. By general theory, the meromorphic connection $(E,\nabla)$ can be associated with the $\mathcal{D}$-module $\mathbb{M}$ over $\mathbb{C}(z)$, introduced in Section 2.1. Identifying $\Omega_{\mathbb{C}(z)/\mathbb{C}}$ with $\mathbb{C}(z)\operatorname{d}\!z$, the map $\nabla:E\rightarrow\Omega_{\mathbb{C}(z)/\mathbb{C}}\otimes E$ provides a differential module structure on $\mathbb{M}$, see [21, Observation 1.8].
@@
because of the residue theorem. According to [15, Theorem 1.3], this minimally twisted local form is equivalent to fixing the irregular parts:
@@
with corresponding residue eigenvalues $0,b_{1},b_{2}$. Here $q_{0,1}$ is a 2-valued function, depending on the values of the square root. The coefficient $\lambda_{2}$ depends only on $b_{0},b_{1}$. See [13, Equations (4.5a), (4.5b)].
@@
The particular cases of interest to us can be calculated directly using the quadratic and cubic formulas. See [15, Theorem 2.1] for a more general statement in arbitrary rank. ∎
@@ 4 Betti spaces
for some $k,l,N\in\mathbb{Z}_{+}$, so that $(\lambda_{l}-\lambda_{l}^{\prime})\neq 0$, and at least one of $\lambda_{k},\lambda_{k}^{\prime}$ is nonzero. In the untwisted case $N=1$, it follows from our assumptions (for instance, that the $a_{i}$’s are different in (JKTVI)) that $k=l$. Let us use the notation $\varepsilon=e^{2\pi\sqrt{-1}/N}$ for a primitive $N$-th root of unity and let $w^{1/N}$ be a fixed sheet of the $N$-valued function on some angular sector of opening $<2\pi$. In the twisted cases $N>1$, and if $q_{i},q_{j}$ belong to the same group of roots under ramification, then for all $1\leq s\leq k$ we have $\lambda_{s}^{\prime}=\lambda_{s}\varepsilon^{-as}$ for some $a\in\mathbb{Z}$. In this case, $k>l$ holds if $N|k$. Recall now from [22, Definition 7.18] that the Stokes directions for the pair $\{q_{i}-q_{j}\}$ form the set of $\varphi\in S^{1}$ that are tangent at 00 to the smooth components of the curve
@@
We refer to [22, Section 7] for the notion of Stokes map $S_{\varphi}\in\textrm{SL}(n,\mathbb{C})$ associated to a connection with irregular singularity at $c$ and Stokes direction $\varphi$. Loosely speaking, $S_{\varphi}$ is the unipotent matrix describing the change of preferred solutions (i.e., ones that have the expected asymptotic expansions) from one side of the Stokes direction to the other. Moreover, let $L_{c}\in\textrm{GL}(n,\mathbb{C})$ be the parallel transport map from a fixed base point $c_{0}$ to $c$ (called the link automorphism). The monodromy automorphism (or topological monodromy) about $c$ is then given by
@@
For the exponential tori in the $JKT*$ cases, see Lemma 4.3. Recall [22, Section 3.2] that the differential Galois group of the differential module corresponding to $\nabla$ over $\mathbb{C}(z)$ is the linear algebraic group generated by the exponential torus and the formal monodromy.
@@ 4.1 The logarithmic case
For the sake of completeness, we mention this case, even though it does not belong to the cases we investigate here. In this case, the connection has three logarithmic singular points on $\mathbb{C}P^{1}$. In our previous paper, we completely described this case from the Betti point of view [7, Section 3]. Here, given that we only have regular singular points, the Stokes phenomenon does not play a role, and the monodromy data can be described by the traces of monodromy matrices on based loops around the three points (which we called Lawton’s trace coordinates [16]). Note that in arbitrary rank, [20, Theorem 1.3] states that traces of products of monodromy matrices generate the invariant ring.
@@ 4.2 The $JKTVI$ case
As we mentioned before, it was proven in [10] that Painlevé VI also arises as isomonodromy deformation of a rank three system of linear differential equations with an irregular singularity of Poincaré rank $1$ and a logarithmic singularity. This case was partially analyzed in the paper [21, Example 2.2], see also [5, Section 11.1]. Here, we describe the Stokes data and the wild Betti space.
@@
For special choice of parameters (say $(\alpha,\beta,\gamma)=(\varepsilon^{2},\varepsilon,1)$), we recover the equation of the Painlevé VI-type Betti moduli space of [21, Subsection 3.1]. More precisely, we get a family of affine cubic surfaces ${\mathcal{M}}_{B}^{JKTVI}\rightarrow\mathbb{C}^{4}$ over the parameter space $\mathcal{P}=\mathbb{C}^{4}$.
@@ 4.3 The $JKTV$ case
This equation is in agreement with the one in [21, Subsection 3.2], for the Painlevé V case. For a suitable choice of $c_{i}$, we can obtain the parameters $s_{i}$ in [21], and ${\mathcal{M}}_{B}^{JKTV}\rightarrow\mathcal{P}=\mathbb{C}\times\mathbb{C}\times\mathbb{C}^{\times}$ is indeed a family of affine cubic surfaces (with our notation the coordinates of the base are $p,q\in\mathbb{C}$ and $\alpha\in\mathbb{C}^{\times}$).
@@ 4.4 The $JKTIVa$ case
This Painlevé IV-type Betti moduli space is a family of affine cubic surfaces over parameter space $\mathcal{P}=\mathbb{C}\times\mathbb{C}$. There is an identification for the parameters $s_{1},s_{2}$ in [21, Subsection 3.7] with the current ones, which we explain here. The equation of the monodromy space for the PIV family in [21] is:
@@ 4.5 The $JKTIVb$ case
Here $s_{1}=\alpha^{1/2}\gamma^{1/2}+\alpha^{-1/2}\gamma^{-1/2}$, $s_{2}=\alpha^{-1/2}\gamma^{1/2}$ provides the exact correspondence with the parameters $s_{1},s_{2}$ in [21, Subsection 3.7].
@@ 4.6 The $JKTII$ case
We obtained the same one-parameter family of affine cubic surfaces ${\mathcal{M}}_{B}^{JKTII}\rightarrow\mathcal{P}=\mathbb{C}^{\times}$ as in [21, Subsection 3.9], up to the change of parameter $\alpha^{-1}\mapsto\alpha$.
@@ 4.7 The $JKTI$ case
This Painlevé I-type Betti moduli space is exactly same as the one in [21, Subsection 3.10].
@@
Table 2: Summary of results on rank three character varieties, cf. the rank two Painlevé cases [21, Table 1.].

### v2

@@ 1 Introduction
Since their discovery at the beginning of the 20th century, the Painlevé equations have attracted a lot of attention, and found a plethora of applications ranging from the theory of random matrices to integrable systems and quantum field theory. A classical result in the field due to Fuchs and Garnier (see Jimbo and Miwa [14, Appendix C]) is that the Painlevé equations all admit descriptions as isomonodromy equations of rank $2$ connections with (regular and irregular) singularities on the Riemann sphere. Meanwhile, Flaschka and Newell [11] found a different such rank $2$ isomonodromy equation description of the second Painlevé equation Painlevé II. This development pointed out that the Painlevé equations may admit several isomonodromic interpretations. Such alternative interpretations (known in the literature as Lax pair representations) are important in order to recognize the relevance of the Painlevé equations for further potential applications. Another Lax pair representation, this time in terms of rank $3$ singular connections, for the Painlevé VI equation was found by Harnad [12, Section 3c]. Subsequently, it was proved by Mazzocco [20] that Harnad’s duality can in fact be considered as an instance of Fourier–Laplace transformation of $\mathcal{D}$-modules. Later, Joshi, Kitaev and Treharne [15][16] gave a list of rank $3$ Lax (or, using their terminology, Fuchs–Garnier) pair representations of all the Painlevé equations. [^3] We note, however, that for Painlevé III they did not quite find a connection, because the differential term is multiplied by a singular matrix. More generally, in [6, Sections 9, 10], Boalch described a way to produce several Lax representations of a class of wild character varieties, namely the ones that are multiplicative quiver varieties with a ”supernova graph” as underlying graph. By assumption, these cases must have unramified irregular type. Among the Painlevé cases studied here, this applies to IVb and VI. Finally, in [8, Section 4] Douçot upgraded Boalch’ quiver variety interpretation to include the rank $2$ and $3$ Lax representations of all Painlevé spaces; notice though that, contrary to the present paper, Douçot works with the spaces of connections, i.e. on the de Rham side.
@@
Character varieties of (irregular) surfaces are of interest and admit applications throughout mathematics and mathematical physics. There are also quite interesting recent conjectures about them, for instance the Geometric P=W conjecture. Even though many theoretical results are known in the irregular case, relatively few concrete examples exist in the literature, especially in the higher (than $2$) rank twisted case. This study illustrates how the general theory of [6] works in practice, by providing a detailed description of a family of rank $3$ examples that is known to be related to the well-known Painlevé family.
@@ 3 Joshi-Kitaev-Treharne systems
in some holomorphic local trivialization of $E$ around each $c_{i}$, where $Q$ is the irregular type from (3), $\Lambda$ is some constant $3\times 3$ matrix (called the residue matrix), and $O(1)$ refers to the holomorphic part. By general theory, the meromorphic connection $(E,\nabla)$ can be associated with the $\mathcal{D}$-module $\mathbb{M}$ over $\mathbb{C}(z)$, introduced in Section 2.1. Identifying $\Omega_{\mathbb{C}(z)/\mathbb{C}}$ with $\mathbb{C}(z)\operatorname{d}\!z$, the map $\nabla:E\rightarrow\Omega_{\mathbb{C}(z)/\mathbb{C}}\otimes E$ provides a differential module structure on $\mathbb{M}$, see [23, Observation 1.8].
@@
because of the residue theorem. According to [17, Theorem 1.3], this minimally twisted local form is equivalent to fixing the irregular parts:
@@
with corresponding residue eigenvalues $0,b_{1},b_{2}$. Here $q_{0,1}$ is a 2-valued function, depending on the values of the square root. The coefficient $\lambda_{2}$ depends only on $b_{0},b_{1}$. See [15, Equations (4.5a), (4.5b)].
@@
The particular cases of interest to us can be calculated directly using the quadratic and cubic formulas. See [17, Theorem 2.1] for a more general statement in arbitrary rank. ∎
@@ 4 Betti spaces
for some $k,l,N\in\mathbb{Z}_{+}$, so that $(\lambda_{l}-\lambda_{l}^{\prime})\neq 0$, and at least one of $\lambda_{k},\lambda_{k}^{\prime}$ is nonzero. In the untwisted case $N=1$, it follows from our assumptions (for instance, that the $a_{i}$’s are different in (JKTVI)) that $k=l$. Let us use the notation $\varepsilon=e^{2\pi\sqrt{-1}/N}$ for a primitive $N$-th root of unity and let $w^{1/N}$ be a fixed sheet of the $N$-valued function on some angular sector of opening $<2\pi$. In the twisted cases $N>1$, and if $q_{i},q_{j}$ belong to the same group of roots under ramification, then for all $1\leq s\leq k$ we have $\lambda_{s}^{\prime}=\lambda_{s}\varepsilon^{-as}$ for some $a\in\mathbb{Z}$. In this case, $k>l$ holds if $N|k$. Recall now from [24, Definition 7.18] that the Stokes directions for the pair $\{q_{i}-q_{j}\}$ form the set of $\varphi\in S^{1}$ that are tangent at 00 to the smooth components of the curve
@@
We refer to [24, Section 7] for the notion of Stokes map $S_{\varphi}\in\textrm{SL}(n,\mathbb{C})$ associated to a connection with irregular singularity at $c$ and Stokes direction $\varphi$. Loosely speaking, $S_{\varphi}$ is the unipotent matrix describing the change of preferred solutions (i.e., ones that have the expected asymptotic expansions) from one side of the Stokes direction to the other. Moreover, let $L_{c}\in\textrm{GL}(n,\mathbb{C})$ be the parallel transport map from a fixed base point $c_{0}$ to $c$ (called the link automorphism). The monodromy automorphism (or topological monodromy) about $c$ is then given by
@@
For the exponential tori in the $JKT*$ cases, see Lemma 4.3. Recall [24, Section 3.2] that the differential Galois group of the differential module corresponding to $\nabla$ over $\mathbb{C}(z)$ is the linear algebraic group generated by the exponential torus and the formal monodromy.
@@ 4.1 The logarithmic case
For the sake of completeness, we mention this case, even though it does not belong to the cases we investigate here. In this case, the connection has three logarithmic singular points on $\mathbb{C}P^{1}$. In our previous paper, we completely described this case from the Betti point of view [9, Section 3]. Here, given that we only have regular singular points, the Stokes phenomenon does not play a role, and the monodromy data can be described by the traces of monodromy matrices on based loops around the three points (which we called Lawton’s trace coordinates [18]). Note that in arbitrary rank, [22, Theorem 1.3] states that traces of products of monodromy matrices generate the invariant ring.
@@ 4.2 The $JKTVI$ case
As we mentioned before, it was proven in [12] that Painlevé VI also arises as isomonodromy deformation of a rank three system of linear differential equations with an irregular singularity of Poincaré rank $1$ and a logarithmic singularity. This case was partially analyzed in the paper [23, Example 2.2], see also [5, Section 11.1]. Here, we describe the Stokes data and the wild Betti space.
@@
For special choice of parameters (say $(\alpha,\beta,\gamma)=(\varepsilon^{2},\varepsilon,1)$), we recover the equation of the Painlevé VI-type Betti moduli space of [23, Subsection 3.1]. More precisely, we get a family of affine cubic surfaces ${\mathcal{M}}_{B}^{JKTVI}\rightarrow\mathbb{C}^{4}$ over the parameter space $\mathcal{P}=\mathbb{C}^{4}$.
@@ 4.3 The $JKTV$ case
This equation is in agreement with the one in [23, Subsection 3.2], for the Painlevé V case. For a suitable choice of $c_{i}$, we can obtain the parameters $s_{i}$ in [23], and ${\mathcal{M}}_{B}^{JKTV}\rightarrow\mathcal{P}=\mathbb{C}\times\mathbb{C}\times\mathbb{C}^{\times}$ is indeed a family of affine cubic surfaces (with our notation the coordinates of the base are $p,q\in\mathbb{C}$ and $\alpha\in\mathbb{C}^{\times}$).
@@ 4.4 The $JKTIVa$ case
This Painlevé IV-type Betti moduli space is a family of affine cubic surfaces over parameter space $\mathcal{P}=\mathbb{C}\times\mathbb{C}$. There is an identification for the parameters $s_{1},s_{2}$ in [23, Subsection 3.7] with the current ones, which we explain here. The equation of the monodromy space for the PIV family in [23] is:
@@ 4.5 The $JKTIVb$ case
Here $s_{1}=\alpha^{1/2}\gamma^{1/2}+\alpha^{-1/2}\gamma^{-1/2}$, $s_{2}=\alpha^{-1/2}\gamma^{1/2}$ provides the exact correspondence with the parameters $s_{1},s_{2}$ in [23, Subsection 3.7].
@@ 4.6 The $JKTII$ case
We obtained the same one-parameter family of affine cubic surfaces ${\mathcal{M}}_{B}^{JKTII}\rightarrow\mathcal{P}=\mathbb{C}^{\times}$ as in [23, Subsection 3.9], up to the change of parameter $\alpha^{-1}\mapsto\alpha$.
@@ 4.7 The $JKTI$ case
This Painlevé I-type Betti moduli space is exactly same as the one in [23, Subsection 3.10].
@@
Table 2: Summary of results on rank three character varieties, cf. the rank two Painlevé cases [23, Table 1.].